UROP Proposal

The Fishwrap Project for Personalized Newspapers

April 27, 1994

I will be continuing work with Pascal Chesnias on the Fishwrap Project, a project at the Media Lab to generate personalized newspapers. This project was started in the Newspapers of the Futures freshman advisor seminar. The project will provide participants with a personalized newspaper including news from one’s hometown and news on topics of special interest.

Specifically, I will be doing work on generating categories for news stories, writing software to parse incoming information from news providers, and work on the interface, specifically, moving a large portion of the interface to a World Wide Web Server. Over the course of the UROP, my tasks may expand to include other parts of the Fishwrap Project.

I have extensive programming experience which will be applied in this project. Additionally, I am one of the maintainers of the software interface that is being used for the Fishwrap Project, and I am very familiar with it and will be able to make enhancements to the interface more easily. This work will be a continuation of this past semesters work.